OpenStudy (anonymous):

An oil company distributes oil in a metal can shaped like a cylinder that has an actual radius of 5.1cm and a height of 15.1cm. A worker incorrectly measured the radius as 5cm and the height as 15cm. Determine the relative error in calculating the surface area, to the nearest thousandth.

OpenStudy (pfenn1):

The formula to figure the surface area of a cylinder is\[SA=2\pi r^2 +2\pi rh\]where r is the radius and h is the height.\[SA_{actual}=2\pi (5.1)^2+2\pi(5.1)(15.1)=\]\[SA_{measured}=2\pi (5)^2+2\pi(5)(15)=\]